OpenStudy (anonymous):

Use the discriminant to determine the nature of the roots of the following equation. x2 + 2x + 5 = 0 Double root real and rational root real and irrational root imaginary root

OpenStudy (unklerhaukus):

A Quadratic Equation\[ax^2+bx+c=0\] The Quadratic Formula\[x_{1,2}=\frac{-b\pm\sqrt{b^2-4ac}}{2a}\] Factored for of the Quadratic Equation\[(x-x_1)(x-x_2)=0\] start by identifying the coefficients a,b,c Compare \[x^2+2x+5=0\] With \[ax^2+bx+c=0\] \(a=?\) \(b=?\) \(c=?\) The radicand of the quadratic formula is called the Discriminant \(\Delta =b^2-4ac\) if \(\Delta=0\) there is only one solution to the quadratic equation if \(\Delta>0\) there are two solutions if \(\Delta <0\) the solutions are complex another way of putting this if \(\Delta=0\) there is two equal solutions if \(\Delta>0\) there are two different solutions if \(\Delta<0\) there are no real solutions if \[\Delta \neq 0= d^2 \] the solutions are rational then taking the square root in the quadratic formula will not leave any irrational component \[x_{1,2}=\frac{-b\pm \sqrt\Delta}{2a}=\frac{-b\pm \sqrt d^2}{2a}=\frac{-b\pm d}{2a}\]

OpenStudy (anonymous):

"No real roots" isn't one of the answer choices.

OpenStudy (unklerhaukus):

What are your values for \(a\), \(b\), \(c\)? What do you get for the discriminant ? \[\Delta=b^2-4ac\]

OpenStudy (anonymous):

Okay, thanks!

OpenStudy (unklerhaukus):

What do you get?

OpenStudy (unklerhaukus):

\[\Delta = (2)^2-4\times(1)\times(5)=\dots\]

OpenStudy (unklerhaukus):

"No real roots" is the same as "complex roots only "